OpenStudy (pulsified333):

Suppose that f(t) is a function for which f(70)=8.3 and f′(70)=5.5. Use this information to estimate the following. Round your answer to one decimal place. 1. f(70+1)≈ 2. f(70−1)≈

OpenStudy (pulsified333):

@UnkleRhaukus

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

if the run is +1, the rise will be about 5.5 if the run is -1, the rise will be about -5.5

OpenStudy (pulsified333):

ok

OpenStudy (unklerhaukus):

what is the best guess for the height of the function at x = 70+1?

OpenStudy (pulsified333):

umm

OpenStudy (pulsified333):

9.3? or 13.8?

OpenStudy (pulsified333):

i dont know

OpenStudy (unklerhaukus):

f(70+1) ≈ f(70) + 1 x f'(70) = 8.3 + 5.5 = 13.8

OpenStudy (unklerhaukus):

f(70-1) ≈ f(70) - 1 x f'(70) = 8.3 - 5.5 =

OpenStudy (pulsified333):

wait so what I did was right?

OpenStudy (unklerhaukus):

yeah 13.8 is the best (linear) approximation for f(71)

OpenStudy (unklerhaukus):

What do you get for f(70-1) ?

OpenStudy (pulsified333):

2.8

OpenStudy (unklerhaukus):

Great work!

OpenStudy (pulsified333):

thanks :D